solve the equation for the positive value of x. Show work.
2 logx - log(x-1) = log4

Answer by lwsshak3(11628)

solve the equation for the positive value of x. Show work.
2 logx - log(x-1) = log4
log[x^2/(x-1)]=log(4)
x^2/(x-1)=4
x^2=4x-4
x^2-4x+4=0
(x-2)^2=0
x=2 (multiplicity 2)